# BAT: Learning Event-based Optical Flow with Bidirectional Adaptive Temporal Correlation

### v1

@@ Abstract
Event cameras deliver visual information characterized by a high dynamic range and high temporal resolution, offering significant advantages in estimating optical flow for complex lighting conditions and fast-moving objects. Current advanced optical flow methods for event cameras largely adopt established image-based frameworks. However, the spatial sparsity of event data limits their performance. In this paper, we present BAT, an innovative framework that estimates event-based optical flow using bidirectional adaptive temporal correlation. BAT includes three novel designs: 1) a bidirectional temporal correlation that transforms bidirectional temporally dense motion cues into spatially dense ones, enabling accurate and spatially dense optical flow estimation; 2) an adaptive temporal sampling strategy for maintaining temporal consistency in correlation; 3) spatially adaptive temporal motion aggregation to efficiently and adaptively aggregate consistent target motion features into adjacent motion features while suppressing inconsistent ones. Our results rank $1^{st}$ on the DSEC-Flow benchmark, outperforming existing state-of-the-art methods by a large margin while also exhibiting sharp edges and high-quality details. Notably, our BAT can accurately predict future optical flow using only past events, significantly outperforming E-RAFT’s warm-start approach. Code: https://github.com/gangweiX/BAT.
1 Introduction
@@
Optical flow has been an enduring fundamental task in the field of computer vision, which can be applied to a broad range of practical tasks, including video reconstruction [37, 38, 45] and visual odometry [14, 48, 52, 30] for robotics. Image-based optical flow estimation [4, 11, 36, 8, 47, 27, 40] has demonstrated remarkable performance on public benchmarks [1, 24, 22]. Nonetheless, these methods struggle with motion blur and over-exposed or under-exposed regions due to the limited dynamic range of conventional image sensors. Event cameras present a promising alternative for optical flow estimation. Compared to image-based cameras, event cameras continuously capture asynchronous brightness changes, providing high temporal resolution without motion blur (microseconds), along with a high dynamic range and low power consumption. Therefore, event-based optical flow estimation is more advantageous for fast-moving objects, high dynamic range scenes, and resource-constrained devices.
@@
Figure 1: Left: Comparison with state-of-the-art event-based flow methods on DSEC-Flow [5] and MVSEC [51] benchmarks. Our BAT achieves the highest accuracy among all published methods. Right: Compared to E-RAFT [6] and TMA [16], our BAT shows much higher efficiency in iterative optimization. Additionally, the aforementioned methods converge at a high error rate quickly due to limited unidirectional motion cues. In contrast, our BAT fully utilizes bidirectional and rich motion cues, allowing for continuous optimization for better results.
@@
However, the data captured by event cameras is spatially sparse, irregular, and asynchronous, making it unsuitable for directly applying current state-of-the-art image-based flow methods to estimate event-based optical flow. To leverage established convolutional neural network architectures [4, 11, 36], recent works [50, 6] transform consecutive event streams into grid-like/tensor representations and input them into these networks to predict the optical flow.
@@
A representative method is E-RAFT [6]. Inspired by RAFT [36], E-RAFT employs image-based optical flow architectures for event-based optical flow estimation. It first converts two event sequences into grid-like/tensor representations, constructs a correlation volume between the two representations, and then regresses the optical flow from the correlation volume. The correlation volume, which measures visual similarity between pixels in two images, has become an essential component in mainstream image-based/frame-based optical flow estimation methods. However, unlike dense and rich RGB or grayscale image data, event data is spatially sparse, resulting in the constructed correlation volume with insufficient motion cues, which significantly limits the performance of this approach. To provide more motion cues, TMA [16] computes the correlation volumes between the reference event frame and multiple forward event frames. Event data is inherently sequential. Although TMA has achieved improved results, it relies solely on forward temporal motion cues and overlooks backward temporal motion cues, which limits its overall performance (see Fig. 1).
@@
Extensive experiments demonstrate that our approach outperforms previous methods by a large margin on public benchmarks. Specifically, our BAT ranks $1^{st}$ on the DSEC-Flow [5] benchmark, exceeding E-RAFT [6] and TMA [16] by 39.45% and 28.98%, respectively, for the 1PE metric.
Additionally, the proposed backward temporal correlation offers two notable advantages: first, it enables us to predict future optical flow using only past events, a capability crucial for interpreting and interacting with dynamic environments, as shown in Fig. 7; second, it effectively handles occlusions caused by objects moving out of the target frame, as shown in Fig. 2.
@@
  • Our BAT ranks $1^{st}$ on the DSEC-Flow benchmark, outperforming existing published methods by a large margin.
@@ 2 Related Work
Deep learning has demonstrated remarkable success in optical flow estimation tasks. The pioneering optical flow method, FlowNet [4], introduced an end-to-end convolutional neural network that directly takes two images as input and outputs optical flow, achieving impressive performance. Subsequently, network architectures [32, 36, 8, 47, 3] and training strategies [40, 34] for optical flow estimation have been continuously improved, leading to consistent advancements in performance. For example, PWC-Net [32, 33] and LiteFlowNet [9, 10] introduced pyramid and warping-based cost volumes to estimate optical flow in a coarse-to-fine manner. These methods significantly reduce inference time while enhancing accuracy. RAFT [36] presents a novel deep network architecture for optical flow, which first builds a 4D correlation volume for all pairs of pixels, and then iteratively regresses the optical flow from it. Combined with its new training strategy, RAFT sets a new state-of-the-art in optical flow estimation.
@@
Inspired by RAFT, several novel designs, such as global attention [12], kernel patch attention [19], and super kernels [35], have been proposed to improve the performance of motion aggregation. Recently, many works [8, 47, 18, 31, 41] have introduced transformers into optical flow estimation to enhance global matching. These methods have all led to improvements in accuracy. Efforts [46, 49, 13, 44] have also been made to overcome the excessive memory consumption introduced by the high-resolution 4D correlation volume in RAFT. Instead of focusing on a two-frame setting, VideoFlow [28] concurrently estimates bidirectional optical flows for multiple consecutive frames. In contrast to VideoFlow, our approach builds on the sparse nature of event data and transforms the rich temporal motion cues into the target frame, focusing on estimating its motion. UnFlow [23] designs an unsupervised loss on bidirectional optical flow computed from two static images. In contrast, our method computes bidirectional temporal correlation for more accurate and detailed forward optical flow prediction.
@@
Learning-based event-based flow methods also largely draw upon well-established image-based optical flow architectures. Moreover, adapting to these architectures often requires converting events into tensor representations, like event frames or voxel grids. These learning-based methods can be categorized into supervised [6, 16, 42], semi-supervised [50, 15, 2], and unsupervised [26, 25, 7, 52]. Unsupervised methods rely solely on event data; semi-supervised methods use grayscale images as a supervisory signal; supervised methods require accurate ground truth flow for supervision. Currently, supervised methods dominate on standard benchmarks [5, 51]. A representative approach is E-RAFT [6], which introduces the RAFT [36] architecture into event-based optical flow estimation. Subsequently, TMA [16] extends the two-frame correlation volume of E-RAFT to a multi-frame correlation volume, providing more temporal motion cues. ADMFlow [20] renders high-quality datasets to facilitate event-flow learning. Without using a correlation volume, IDNet [42] introduces iterative deblurring for accurate event-based flow estimation.
@@ 3 Method
Figure 3: Network architecture of the proposed BAT. We first split the reference and target event streams into multiple groups and extract the corresponding features separately. Then, we perform forward and backward temporal correlations between the sequential event features. We further propose spatially adaptive temporal motion aggregation, which integrates temporally consistent target motion features into adjacent motion features while suppressing inconsistent ones. Finally, the optical flow is iteratively updated through a ConvGRU.
@@
Building on the spatial sparsity and temporal density of event data, we propose BAT, a novel framework that transforms temporally dense motion cues into spatially dense ones, enabling spatially dense and accurate optical flow estimation. The overview of the proposed network framework is illustrated in Figure Fig. 3. Before feeding the event data into the network, the event stream is first converted into an event voxel grid representation. In this section, we first describe the event voxel grid representation and then detail the network architecture.
3.1 Event Representation
@@
To ensure compatibility with CNNs and facilitate correlation volume construction, we follow previous works [6, 16] by transforming the event stream $\mathcal{E}(t_{i},t_{i+1})$ into an event voxel grid $\bm{V}\in\mathbb{R}^{B\times H_{0}\times W_{0}}$ with discretized time dimension $B$. Specifically, given an event stream $\mathcal{E}(t_{i},t_{i+1})$, $\bm{V}(b,x,y)$ is obtained by:
@@
3.2 Feature Extraction
Given two consecutive event streams $\mathcal{E}(t_{i-1},t_{i})$ and $\mathcal{E}(t_{i},t_{i+1})$, where $t_{i+1}-t_{i}=t_{i}-t_{i-1}=\Delta t$, our goal is to estimate the optical flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$ form timestamp $t_{i}$ to $t_{i+1}$. Firstly, we convert the event streams into voxel grid representations, denoted as $\bm{V}_{0}$ and $\bm{V}_{1}$, respectively, using Eq. 1.
To exploit temporally dense motion cues, we split both $\bm{V}_{0}$ and $\bm{V}_{1}$ into $N$ groups equally along the time bin dimension, and each group contains $B/N$ time bins. Then we input the voxel grid event representations of these 2N groups into a shared-weight feature extraction network to produce a series of features $\bm{F}_{n}$ ($\bm{F}_{n}\in\mathbb{R}^{D\times H\times W}$), $n=1,2,\dots,2N$, where $D$, $H$ and $W$ denote the channel, height, and width of feature $\bm{F}_{n}$. The feature extraction network consists of 6 residual blocks, similar to RAFT [36]. Among these temporal features, $\bm{F}_{N}$ serves as the reference frame, while $\bm{F}_{2N}$ serves as the target frame.
Figure 4: Spatially Adaptive Temporal Motion Aggregation (SATMA).
3.3 Bidirectional Temporal Correlation
Correlation volume, which measures pixel-wise visual similarities between frames, provides crucial information for flow estimation. However, existing flow methods often focus on the two-frame setting [6] or unidirectional correlation [16]. In contrast, we propose bidirectional temporal correlation (BTC) to transform temporally rich motion cues into spatially dense ones. BTC consists of forward temporal correlation and backward temporal correlation. Forward temporal correlation computes the correlation between the reference frame $\bm{F}_{N}$ and its next frames $\bm{F}_{N+j}$ ($j=1,\dots,N$), and backward temporal correlation computes the correlation between the reference frame $\bm{F}_{N}$ and its previous frames $\bm{F}_{N-j}$ ($j=1,2,\dots,N-1$).
@@ 3.3 Bidirectional Temporal Correlation
${\bm{C}}_{j}^{fwd}=\frac{{\bm{F}}_{N}\bm{\cdot}\mathcal{W}({\bm{F}}_{N+j};j\bm{df})}{\sqrt{D}},j=1,2,\dots,N,$ (3)
where $\bm{\cdot}$ denotes the dot product, $\mathcal{W}(\cdot;\cdot)$ denotes the warping of adjacent frames to the reference frame using optical flow. Specifically, for each pixel $p$ in $\bm{F}_{N}$, we use the current flow $j\bm{df}$ to find its corresponding point $p^{\prime}$ in $\bm{F}_{N+j}$, then define a local grid around $p^{\prime}$ within a radius $r$ and bilinearly sample the values from the local grid. The number of samples for the local grid is $(2r+1)^{2}$.
@@
$\bm{C}_{j}^{bwd}=\frac{\bm{F}_{N}\bm{\cdot}\mathcal{W}(\bm{F}_{N-j};(-j)\bm{df})}{\sqrt{D}},j=1,2,\dots,N-1.$ (4)
Finally, we obtain $N$ groups of forward temporal correlation maps and $N\!-\!1$ groups of backward temporal correlation maps. Through this bidirectional temporal correlation, we transform temporally dense motion cues into spatially dense ones, enabling spatially detailed and accurate optical flow estimation. In particular, the backward temporal correlation maps are highly beneficial for handling occlusions of objects moving out of the target frame, as shown in Fig. 2.
Adaptive Temporal Sampling Strategy. The sampling radius $r$ is set as a hyperparameter in previous works [6, 16]. However, as motion patterns change over time, a manually chosen $r$ may not be robust for temporal consistent correlation. To address this, we introduce a learnable parameter $\alpha$ as a scale factor to obtain a learnable sampling radius $lr$,
@@
3.4 Spatially Adaptive Temporal Motion Aggregation
Effectively utilizing temporally rich motion cues is crucial for achieving spatially dense and accurate optical flow estimation. However, due to the non-uniformity of optical flow over time, linearly computing correlations on adjacent temporal frames introduces motion features inconsistent with the target frame. To address this, we propose a spatially adaptive temporal motion aggregation (SATMA) module that adaptively enhances motion features consistent with the target frame while suppressing inconsistent ones.
@@ 3.4 Spatially Adaptive Temporal Motion Aggregation
Fig. 4 shows the overview of the spatially adaptive temporal motion aggregation module. Given the target motion feature $\bm{M}_{N}^{fwd}$ and an adjacent motion feature $\bm{M}_{j}^{fwd}$ (or $\bm{M}_{j}^{bwd}$), we first concatenate them, then pass them through a convolutional layer followed by a sigmoid activation function to generate a spatial attention map $\bm{A}_{spa}$. The spatial attention effectively enhances consistent motion features while suppressing inconsistent ones. We also adopt deformable attention [43] to efficiently integrate consistent or similar motion features from $\bm{M}_{N}^{fwd}$ into $\bm{M}_{j}^{fwd}$. Specifically, $\bm{M}_{j}^{fwd}$ is first projected linearly to obtain the query $\bm{Q}$, which is then fed into a lightweight offset network to generate sparse sampling locations. Based on these locations, we can sample $\bm{M}_{N}^{fwd}$ to obtain the key $\bm{K}$ and value $\bm{V}$. Then we obtain the aggregated motion feature $\bm{M}^{agg}_{j}$ through attention:
@@
MB MultiCM [29] 3.472 76.570 48.480 30.855 13.983
SL E-RAFT [6] 0.788 12.742 4.740 2.684 2.851
ADMFlow [20] 0.779 12.522 4.673 2.647 2.838
EEMFlow+ [21] 0.751 11.403 3.932 2.145 2.669
TMA [16] 0.743 10.863 3.972 2.301 2.684
IDNet [42] 0.719 10.069 3.497 2.036 2.723
@@
Table 1: DSEC [5] benchmark evaluation. Our BAT ranks $1^{st}$ among all published methods. MB refers to model-based methods that do not require training data, and SL represents supervised learning methods that require ground truth optical flow. Bold: best.
3.5 Loss Function
Following RAFT [36], we supervised our network using the $l_{1}$ loss between the predicted and ground truth flow over the entire sequence of predictions, $\bm{f}^{1},\bm{f}^{2},\cdots,\bm{f}^{K}$}, applying exponentially increasing weights. Given the ground truth flow $\bm{f}^{gt}$, the loss is defined as,
$\mathcal{\bm{L}}=\sum_{i=1}^{K}\gamma^{K-i}||{\bm{f}}^{i}-{\bm{f}}^{gt}||_{1},$ (9)
@@ 4 Experiments
[…]
MB MultiCM [29] 0.30 0.10 1.25 9.21
SSL EV-FlowNet [50] 0.49 0.20 1.23 7.30
Spike-FlowNet [15] 0.49 - 1.09 -
STE-FlowNet [2] 0.42 0.00 0.99 3.90
SL E-RAFT [6] 0.27 0.00 0.84 1.70
DCEIFlow [39] 0.91 0.71 1.87 19.1
ADMFlow [20] 0.52 0.00 1.91 19.2
TMA [16] 0.25 0.07 0.70 1.08
IDNet [42] 0.29 0.00 0.75 1.20
@@
Table 2: MVSEC [51] benchmark evaluation. SSL denotes semi-supervised learning methods that use grayscale images for supervision.
@@
Full model (BAT) 0.655 7.715 2.896 1.773 2.439
Table 3: Ablation study on DSEC-Flow benchmark. BTC represents bidirectional temporal correlation, ATS represents adaptive temporal sampling, and SATMA represents spatially adaptive temporal motion aggregation.
@@
Tgt.+Ref. event MultiCM [29] 3.472 76.570
TamingCM [26] 2.330 68.293
EV-FlowNet [50] 2.320 55.400
@@ 4.1 Experimental Setup
Datasets and evaluation setups. Following previous work, we conduct extensive experiments on two popular event-based datasets: DSEC-Flow [5] and MVSEC [51]. For DSEC-Flow, we train the models on the official training dataset and then evaluate them on the public benchmark. DSEC-Flow provides 8,170 training samples and 416 testing samples, with a resolution of $640\times 480$. For MVSEC, we perform training and evaluation on two types of event data, $dt=4$ and $dt=1$, which are divided based on time intervals.
Evaluation metrics. The end-point error (EPE) is the primary metric used to evaluate the accuracy of optical flow predictions for both DSEC-Flow [5] and MVSEC [51]. Additionally, we report the percentage of ground truth pixels with an optical flow magnitude error greater than N (NPE), where N is either 1, 2, or 3, as well as the Angular Error (AE) for DSEC-Flow. For MVSEC, we report the percentage of ground truth pixels with EPE exceeding 3 and 5% of the flow magnitude (%Out).
@@
Implementation details. We implement our model with PyTorch and perform our experiments using NVIDIA RTX 3090 GPUs. We set the temporal group $N$ to 3 in the final BAT model. For experiments on DSEC-Flow, we set the event representation to $B=15$ time bins for every 100ms of events. For MVSEC, $B$ is set to 5 when $dt=1$ and 15 when $dt=4$. The length $B$ of time bins is kept consistent with E-RAFT [6] and TMA [16]. The final channel dimension $D$ for feature extraction is 128, and the flow updater is identical to that of RAFT [36]. The sampling radius $r$ is set to 2. For all training, we use the AdamW [17] optimizer with a one-cycle learning rate schedule, setting the learning rate to 2e-4. We use 8 update iterations during training, and the weight $\gamma$ is set to 0.8. For DSEC-Flow, we train our BAT for 200k steps with a batch size of 8, while for MVSEC, we train it for 100k steps.

### v2

@@ Abstract
Event cameras deliver visual information characterized by a high dynamic range and high temporal resolution, offering significant advantages in estimating optical flow for complex lighting conditions and fast-moving objects. Current advanced optical flow methods for event cameras largely adopt established image-based frameworks. However, the spatial sparsity of event data limits their performance. In this paper, we present BAT, an innovative framework that estimates event-based optical flow using bidirectional adaptive temporal correlation. BAT includes three novel designs: 1) a bidirectional temporal correlation that transforms bidirectional temporally dense motion cues into spatially dense ones, enabling accurate and spatially dense optical flow estimation; 2) an adaptive temporal sampling strategy for maintaining temporal consistency in correlation; 3) spatially adaptive temporal motion aggregation to efficiently and adaptively aggregate consistent target motion features into adjacent motion features while suppressing inconsistent ones. Our BAT achieves state-of-the-art performance on the DSEC-Flow benchmark, outperforming existing methods by a large margin while also exhibiting sharp edges and high-quality details. Our BAT can accurately predict future optical flow using only past events, significantly outperforming E-RAFT’s warm-start approach.
Codehttps://github.com/gangweix/BAT
Introduction
@@
Optical flow has been an enduring fundamental task in the field of computer vision, which can be applied to a broad range of practical tasks, including video reconstruction (tulyakov2021time; tulyakov2022time; xu2024hdrflow) and visual odometry (zhang2024leveraging; shiba2024secrets), and robotics (xu2025pixel; liang2025parameter; liang2025sood++; li2025recogdrive; lin2025prompting). Image-based optical flow estimation (flownet; raft; sea-raft) has demonstrated remarkable performance on public benchmarks (sintel; kitti15; spring). Nonetheless, these methods struggle with motion blur and over-exposed or under-exposed regions due to the limited dynamic range of conventional image sensors. Event cameras present a promising alternative for optical flow estimation. Compared to image-based cameras, event cameras continuously capture asynchronous brightness changes, providing high temporal resolution without motion blur (microseconds), along with a high dynamic range and low power consumption. Therefore, event-based optical flow estimation is more advantageous for fast-moving objects, high dynamic range scenes, and resource-constrained devices.
@@
Figure 1: Left: Comparison with state-of-the-art event-based flow methods on DSEC-Flow and MVSEC benchmarks. Our BAT achieves the highest accuracy among all published methods. Right: Compared to E-RAFT and TMA, our BAT shows much higher efficiency in iterative optimization. Additionally, the aforementioned methods converge at a high error rate quickly due to limited unidirectional motion cues. In contrast, our BAT fully utilizes bidirectional and rich motion cues, allowing for continuous optimization for better results.
@@
However, the data captured by event cameras is spatially sparse, irregular, and asynchronous, making it unsuitable for directly applying current state-of-the-art image-based flow methods to estimate event-based optical flow. To leverage established convolutional neural network architectures (flownet; flownet2; raft), recent works (evflownet; eraft) transform consecutive event streams into grid-like/tensor representations and input them into these networks to predict the optical flow.
@@
A representative method is E-RAFT (eraft). Inspired by RAFT (raft), E-RAFT employs image-based optical flow architectures for event-based optical flow estimation. It first converts two event sequences into grid-like/tensor representations, constructs a correlation volume between the two representations, and then regresses the optical flow from the correlation volume. The correlation volume, which measures visual similarity between pixels in two images, has become an essential component in mainstream image-based/frame-based optical flow estimation methods. However, unlike dense and rich RGB or grayscale image data, event data is spatially sparse, resulting in the constructed correlation volume with insufficient motion cues, which significantly limits the performance of this approach. To provide more motion cues, TMA (tma) computes the correlation volumes between the reference event frame and multiple forward event frames. Event data is inherently sequential. Although TMA has achieved improved results, it relies solely on forward temporal motion cues and overlooks backward temporal motion cues, which limits its overall performance (see Figure 1).
@@
Extensive experiments demonstrate that our approach outperforms previous methods by a large margin on public benchmarks. Specifically, our BAT ranks $1^{st}$ on the DSEC-Flow (dsec) benchmark, exceeding E-RAFT (eraft) and TMA (tma) by 39.45% and 28.98%, respectively, for the 1PE metric.
Additionally, the proposed backward temporal correlation offers two notable advantages: first, it enables us to predict future optical flow using only past events, a capability crucial for interpreting and interacting with dynamic environments, as shown in Figure 6; second, it effectively handles occlusions caused by objects moving out of the target frame, as shown in Figure 2.
@@
  • Our BAT achieves state-of-the-art performance on the DSEC-Flow benchmark, outperforming all previously published methods.
@@
Deep learning has demonstrated remarkable success in image matching tasks (igev; igev++; coatrsnet; monster; zerostereo; fast-acv; cheng2024adaptive). The pioneering flow method, FlowNet (flownet), introduced an end-to-end convolutional neural network that directly takes two images as input and outputs optical flow, achieving impressive performance. Subsequently, network architectures (raft; flowformer) and training strategies (sea-raft; sun2022disentangling) for optical flow estimation have been continuously improved, leading to consistent advancements in performance. For example, PWC-Net (pwcnet; sun2019models) and LiteFlowNet (liteflownet; lightweight2) introduced pyramid and warping-based cost volumes to estimate optical flow in a coarse-to-fine manner. These methods significantly reduce inference time while enhancing accuracy. RAFT presents a novel deep network architecture for optical flow, which first builds a 4D correlation volume for all pairs of pixels, and then iteratively regresses the optical flow from it. Combined with its new training strategy, RAFT sets a new state-of-the-art in optical flow estimation.
@@
Inspired by RAFT, several novel designs, such as global attention (gma), kernel patch attention (kpaflow), and super kernels (skflow), have been proposed to improve the performance of motion aggregation. Recently, many works (flowformer; gmflow; transflow; craft; croco) have introduced transformers into optical flow estimation to enhance global matching. These methods have all led to improvements in accuracy. Efforts (flow1d; hcv; scv; meflow) have also been made to overcome the excessive memory consumption introduced by the high-resolution 4D correlation volume in RAFT. Instead of focusing on a two-frame setting, VideoFlow (videoflow) concurrently estimates bidirectional optical flows for multiple consecutive frames. In contrast to VideoFlow, our approach builds on the sparse nature of event data and transforms the rich temporal motion cues into the target frame, focusing on estimating its motion. UnFlow (unflow) designs an unsupervised loss on bidirectional optical flow computed from two static images. In contrast, our method computes bidirectional temporal correlation for more accurate and detailed forward optical flow prediction.
@@
Learning-based event-based flow methods also largely draw upon well-established image-based optical flow architectures. Moreover, adapting to these architectures often requires converting events into tensor representations, like event frames or voxel grids. These learning-based methods can be categorized into supervised (eraft; tma; idnet), semi-supervised (evflownet; spikeflownet; steflownet), and unsupervised (tamingcm; paredes2021back; hagenaars2021self; zhu2019unsupervised). Unsupervised methods rely solely on event data; semi-supervised methods use grayscale images as a supervisory signal; supervised methods require accurate ground truth flow for supervision. Currently, supervised methods dominate on standard benchmarks (dsec; mvsec). A representative approach is E-RAFT (eraft), which introduces the RAFT (raft) architecture into event-based optical flow estimation. Subsequently, TMA (tma) extends the two-frame correlation volume of E-RAFT to a multi-frame correlation volume, providing more temporal motion cues. ADMFlow (admflow) renders high-quality datasets to facilitate event-flow learning. Without using a correlation volume, IDNet (idnet) introduces iterative deblurring for accurate event-based flow estimation.
@@
Figure 3: Network architecture of the proposed BAT. We first split the reference and target event streams into multiple groups and extract the corresponding features separately. Then, we perform forward and backward temporal correlations between the sequential event features. We further propose spatially adaptive temporal motion aggregation, which integrates temporally consistent target motion features into adjacent motion features while suppressing inconsistent ones.
@@
Building on the spatial sparsity and temporal density of event data, we propose BAT, a novel framework that transforms temporally dense motion cues into spatially dense ones, enabling spatially dense and accurate optical flow estimation. The overview of the proposed network framework is illustrated in Figure 3. Before feeding the event data into the network, the event stream is first converted into an event voxel grid representation. In this section, we first describe the event voxel grid representation and then detail the network architecture.
Event Representation
@@
To ensure compatibility with CNNs and facilitate correlation volume construction, we follow previous works (eraft; tma) by transforming the event stream $\mathcal{E}(t_{i},t_{i+1})$ into an event voxel grid $\bm{V}\in\mathbb{R}^{B\times H_{0}\times W_{0}}$ with discretized time dimension $B$. Specifically, given an event stream $\mathcal{E}(t_{i},t_{i+1})$, $\bm{V}(b,x,y)$ is obtained by:
@@
Feature Extraction
Given two consecutive event streams $\mathcal{E}(t_{i-1},t_{i})$ and $\mathcal{E}(t_{i},t_{i+1})$, where $t_{i+1}-t_{i}=t_{i}-t_{i-1}=\Delta t$, our goal is to estimate the optical flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$ form timestamp $t_{i}$ to $t_{i+1}$. Firstly, we convert the event streams into voxel grid representations, denoted as $\bm{V}_{0}$ and $\bm{V}_{1}$, respectively, using Equ. 1.
To exploit temporally dense motion cues, we split both $\bm{V}_{0}$ and $\bm{V}_{1}$ into $N$ groups equally along the time bin dimension, and each group contains $B/N$ time bins. Then we input the voxel grid event representations of these 2N groups into a shared-weight feature extraction network to produce a series of features $\bm{F}_{n}$ ($\bm{F}_{n}\in\mathbb{R}^{D\times H\times W}$), $n=1,2,\dots,2N$, where $D$, $H$ and $W$ denote the channel, height, and width of feature $\bm{F}_{n}$. The feature extraction network consists of 6 residual blocks, similar to RAFT (raft). Among these temporal features, $\bm{F}_{N}$ serves as the reference frame, while $\bm{F}_{2N}$ serves as the target frame.
Bidirectional Temporal Correlation
Correlation volume, which measures pixel-wise visual similarities between frames, provides crucial information for flow estimation. However, existing flow methods often focus on the two-frame setting (eraft) or unidirectional correlation (tma). In contrast, we propose bidirectional temporal correlation (BTC) to transform temporally rich motion cues into spatially dense ones. BTC consists of forward temporal correlation and backward temporal correlation. Forward temporal correlation computes the correlation between the reference frame $\bm{F}_{N}$ and its next frames $\bm{F}_{N+j}$ ($j=1,\dots,N$), and backward temporal correlation computes the correlation between the reference frame $\bm{F}_{N}$ and its previous frames $\bm{F}_{N-j}$ ($j=1,2,\dots,N-1$).
@@
${\bm{C}}_{j}^{fwd}=\frac{{\bm{F}}_{N}\boldsymbol{\cdot}\mathcal{W}({\bm{F}}_{N+j};j\bm{df})}{\sqrt{D}},j=1,2,\dots,N,$ (3)
[…]
@@
$\bm{C}_{j}^{bwd}=\frac{\bm{F}_{N}\boldsymbol{\cdot}\mathcal{W}(\bm{F}_{N-j};(-j)\bm{df})}{\sqrt{D}},j=1,2,\dots,N-1.$ (4)
Finally, we obtain $N$ groups of forward temporal correlation maps and N1N\!-\!1 groups of backward temporal correlation maps. Through this bidirectional temporal correlation, we transform temporally dense motion cues into spatially dense ones, enabling spatially detailed and accurate optical flow estimation. In particular, the backward temporal correlation maps are highly beneficial for handling occlusions of objects moving out of the target frame, as shown in Figure 2.
Adaptive Temporal Sampling Strategy. The sampling radius $r$ is set as a hyperparameter in previous works (eraft; tma). However, as motion patterns change over time, a manually chosen $r$ may not be robust for temporal consistent correlation. To address this, we introduce a learnable parameter $\alpha$ as a scale factor to obtain a learnable sampling radius $lr$,
@@
Spatially Adaptive Temporal Motion Aggregation
Effectively utilizing temporally rich motion cues is crucial for achieving spatially dense and accurate optical flow estimation. However, due to the non-uniformity of optical flow over time, linearly computing correlations on adjacent temporal frames introduces motion features inconsistent with the target frame. To address this, we propose a spatially adaptive temporal motion aggregation (SATMA) module that adaptively enhances motion features consistent with the target frame while suppressing inconsistent ones. The architecture of SATMA is illustrated in the supplementary material.
@@
The architecture of the spatially adaptive temporal motion aggregation (SATMA) module is presented in the supplementary material. Given the target motion feature $\bm{M}_{N}^{fwd}$ and an adjacent motion feature $\bm{M}_{j}^{fwd}$ (or $\bm{M}_{j}^{bwd}$), we first concatenate them, then pass them through a convolutional layer followed by a sigmoid activation function to generate a spatial attention map $\bm{A}_{spa}$. The spatial attention effectively enhances consistent motion features while suppressing inconsistent ones. We also adopt deformable attention (deform_attn) to efficiently integrate consistent or similar motion features from $\bm{M}_{N}^{fwd}$ into $\bm{M}_{j}^{fwd}$. Specifically, $\bm{M}_{j}^{fwd}$ is first projected linearly to obtain the query $\bm{Q}$, which is then fed into a lightweight offset network to generate sparse sampling locations. Based on these locations, we can sample $\bm{M}_{N}^{fwd}$ to obtain the key $\bm{K}$ and value $\bm{V}$. Then we obtain the aggregated motion feature $\bm{M}^{agg}_{j}$ through attention:
@@
MB MultiCM 3.472 76.570 48.480 30.855 13.983
SL E-RAFT 0.788 12.742 4.740 2.684 2.851
ADMFlow 0.779 12.522 4.673 2.647 2.838
EEMFlow+ 0.751 11.403 3.932 2.145 2.669
TMA 0.743 10.863 3.972 2.301 2.684
IDNet 0.719 10.069 3.497 2.036 2.723
@@
Table 1: DSEC (dsec) benchmark evaluation. MB refers to model-based methods that do not require training data, and SL represents supervised learning methods that require ground truth optical flow. Bold: best.
Loss Function
Following RAFT, we supervised our network using the $l_{1}$ loss between the predicted and ground truth flow over the entire sequence of predictions, $\bm{f}^{1},\bm{f}^{2},\cdots,\bm{f}^{K}$}, applying exponentially increasing weights. Given the ground truth flow $\bm{f}^{gt}$, the loss is defined as,
[…]
@@
$downarrow EPE\downarrow Out\downarrow$
MB MultiCM 0.30 0.10 1.25 9.21
SSL EV-FlowNet 0.49 0.20 1.23 7.30
Spike-FlowNet 0.49 - 1.09 -
STE-FlowNet 0.42 0.00 0.99 3.90
SL E-RAFT 0.27 0.00 0.84 1.70
DCEIFlow 0.91 0.71 1.87 19.1
ADMFlow 0.52 0.00 1.91 19.2
TMA 0.25 0.07 0.70 1.08
IDNet 0.29 0.00 0.75 1.20
@@
Table 2: MVSEC (mvsec) benchmark evaluation. SSL denotes semi-supervised learning methods that use grayscale images for supervision.
@@
Full (BAT) 0.655 7.715 2.896 1.773 2.439
Table 3: Ablation study on DSEC-Flow benchmark. BTC represents bidirectional temporal correlation, ATS represents adaptive temporal sampling. Full model includes SATMA, denoting spatially adaptive temporal motion aggregation.
@@
Tgt.+Ref. event MultiCM 3.472 76.570
TamingCM 2.330 68.293
EV-FlowNet 2.320 55.400
@@
Datasets and evaluation setups. Following previous work, we conduct extensive experiments on two popular event-based datasets: DSEC-Flow (dsec) and MVSEC (mvsec). For DSEC-Flow, we train the models on the official training dataset and then evaluate them on the public benchmark. DSEC-Flow provides 8,170 training samples and 416 testing samples, with a resolution of $640\times 480$. For MVSEC, we perform training and evaluation on two types of event data, $dt=4$ and $dt=1$, which are divided based on time intervals.
@@
Implementation details. We implement our model with PyTorch and perform our experiments using NVIDIA RTX 3090 GPUs. We set the temporal group $N$ to 3 in the final BAT model. For experiments on DSEC-Flow, we set the event representation to $B=15$ time bins for every 100ms of events. For MVSEC, $B$ is set to 5 when $dt=1$ and 15 when $dt=4$. The length $B$ of time bins is kept consistent with E-RAFT and TMA. The final channel dimension $D$ for feature extraction is 128, and the flow updater is identical to that of RAFT. The sampling radius $r$ is set to 2. For all training, we use the AdamW optimizer with a one-cycle learning rate schedule, setting the learning rate to 2e-4. We use 8 update iterations during training, and the weight $\gamma$ is set to 0.8. For DSEC-Flow, we train our BAT for 200k steps with a batch size of 8, while for MVSEC, we train it for 100k steps.